how do i solve 4x^2+4x-15=0
Found 3 solutions by Alan3354, scott8148, checkley77:
Answer by Alan3354(69443)

how do i solve 4x^2+4x-15=0
-----------
There are several methods. Factoring, completing the square, and the quadratic equation.
--------------
You can use Excel, too, and look for zeroes between -15 and +15.

Answer by scott8148(6628)

factoring ___ (2x + 5) (2x - 3) = 0

2x + 5 = 0 ___ x = -5/2

2x - 3 = 0 ___ x = 3/2

Answer by checkley77(12844)

The factors of the a term(4x) are: 1*4 & 2*2
The factors of the term(-15) are -1*15, 1*-15, -3*5, 3*-5
Now the secret is to find the sum of the products that = 4
1*-1+4*15=-1+60=59 reject
1*15+4*-1=15-4=11 reject
1*1+4*-15=1-60=-59 reject
1*-15+4*4=-15+15=31 reject
1*-3+4*5=-3+20=17 reject
1*5+4*-3=5-12=-7 reject
1*3+4*-5=3-20=-17 reject
2*-1+2*15=-2+30=28 reject
************************************
2*-3+2*5=-6+10=4 eureka.
Proof:
Using the FOIL system.
2x*2x+2x*5-3x*2-3*5
4x*2+10x-6x-15
[4x^2+4x-15]
*************************************
2*3+2*-5=6-10-4 reject
------------------------------
4x^2+4x-15=0
(2x-3)(2x+5)=0
2x-3=0
2x=3
x=3/2
x=1.5 ans.
2x+5=0
2x=-5
x=-5/2
x=-2.5 ans.